### v1

@@ 4.4 Limitations
As noted above, this study attempted to compare the ADS crashed vehicle rate to the most comparable benchmark available in the literature. Although these benchmarks could adjust for some factors that are related to crashed vehicle rate (e.g., geographical area, types of roads, vehicle type), other pertinent factors were not controlled for. These include time of day adjustments and adjustments for the driving density within the geographic areas. As the Waymo ADS is being used as a ride-hailing service, the routes the ADS drives are likely differently dispersed than the general driving population and may be biased toward more densely populated areas. There are challenges when matching benchmark crash and VMT data sources that can be adjusted by these parameters. Even if the origin and destinations of the ADS ride-hailing trips match closely to a comparable human driving population, there still may be differences in the route selection between the ADS vehicle and human drivers that affects crash rates. An aggregate-level analysis comparing the ADS to the current crash status quo should include these route choices if the research question is to determine the safety impact of replacing human driving with ADS vehicle driving. It may also be of interest, however, to compare the ADS vehicle crashed vehicle rate compared to human drivers driving the same routes as the ADS vehicle. This latter comparison is closer to an event-level analysis (Favaro et al., 2023a) than an aggregate-level analysis that is the purpose of the current study and different methods than an outcome-based approach may be more effective in this case.
@@
An ADS vehicle operating in a ride hailing context is not always occupied by a human (e.g., before or in between picking up passengers). Therefore, when examining injury outcomes, if the ADS vehicle is unoccupied, there would be a reduced risk of injury for a similar human-only crash. More broadly, risk of injury in a motor vehicle crash can be affected by many factors, such as seat belt use and vehicle model year (as a proxy for available safety systems). A risk-based approach that uses an injury risk function to compute probability of an injury outcome given some set of inputs can be advantageous in normalizing risk in both the benchmark and ADS crashes for comparison, as has been traditionally done in prospective safety benefit analyses (for example, Kusano and Gabler (2012); Scanlon et al. (2017)).
The Waymo RO service studied in this paper was operating over multiple years on different vehicle platforms and with evolving software versions. Just as the human driving population has changing characteristics over time, certain aspects of the Waymo RO ride-hailing service have changed. The operating environment and territory has expanded during this time period. Starting in the area around Chandler, Arizona in 2020, today the Waymo RO ride-hailing service operates in an area that includes multiple cities in metropolitan Phoenix and over the entire area of San Francisco. A limitation of this study is that the entire history of the Waymo RO crash record was compared to human crash data over different time periods. As the results of this study show, several millions of miles are needed to detect differences between an ADS fleet and benchmarks. As ADS fleets continue to increase the number of VMT, additional analyses that are restricted to certain periods of time or platforms could be enabled. One difficulty in comparing ADS and human crash data is that ADS data is published at a much faster cadence than human crash data. The NHTSA SGO crash data is submitted every month and published shortly after. Complete human crash databases generally are published by the calendar year and can take up to 2 years after the completion of the calendar year to be published.
5 Conclusions
This study compared the crashed vehicle rate of the Waymo RO service to human benchmarks for any property damage or injury, police-reported, and any-injury-reported outcomes. The results show that the Waymo RO service has a lower police-reported and any-injury-reported crashed vehicle rate compared to the benchmarks. When considering all locations together, the any-injury crashed vehicle rate was 0.41 incidents per million miles (IPMM) for the ADS vs 2.78 IPMM for the human benchmark, an 85% reduction or a 6.8 times lower rate. Police-reported crashed vehicle rates for all locations together were 2.1 IPMM for the ADS vs. 4.85 IPMM for the human benchmark, a 57% reduction or 2.3 times lower rate. Because the police-reported and any-injury-reported crashed ADS vehicle rates were much lower than the benchmark rates, the ADS any-injury-reported and police-report crashed vehicle rates were statistically significant over a total of 7.14 million RO miles when compared individually in Phoenix and San Francisco and when compared overall. The comparison for Los Angeles was not statistically significant for zero observed ADS events but few miles. The results are more difficult to interpret for the any property damage or injury outcome comparison. In general, the ADS had a lower crashed vehicle rate than the benchmarks. The benchmark rates themselves, however, varied considerably between locations and within the same location. This raises questions whether the benchmark data sources have comparable reporting thresholds (surveillance bias) or if other factors that were not controlled for in the benchmarks (time of day, mix of driving) is affecting the benchmark rates. We assume that the police-reported and any-injury-reported rates are not as affected by this surveillance bias because of a higher lower threshold for reporting. This is the first study to compare overall crashed vehicle rates of ADS RO data only (as opposed to a mix of ADS testing with a human behind the wheel) to human benchmarks that also corrected for biases reported in the literature. The comparison of crash rates to benchmarks are part of the safety determination lifecycle, which itself is a part of a broader ADS safety case. The results show that the Waymo ADS operating in RO configuration has a lower crashed vehicle rate than human drivers when considering police-reported and any-injury-reported crashes. This result provides directional confirmation of the predictions and analyses used prior to deploying RO operations.
[…]
Appendix A Appendix
The table below lists all the crashes included in this analysis. The SGO report ID can be used to look up further details of these events that are published by NHTSA. The “Waymo 1M Mile RO Event #” column lists the integer event identifier that was reported in Victor et al. (2023) that described all contact events (regardless of whether they were reported as part of the NHTSA SGO). Of the 20 events reported in Victor et al. (2023), 11 were reported as part of the NHTSA SGO. Waymo started to operate in RO before the NHTSA SGO was enacted in July 2021 and as a result, 4 contact incidents that were previously reported in Victor et al. (2023) were not subject to the SGO reporting requirements. The authors retroactively reviewed these 4 contact incidents and determined that 2 of the contact incidents would have qualified for NHTSA SGO reporting. These two (2) incidents were included in the data analyzed for this paper. The location column refers to the market the crash occurred in (PHX for Phoenix, SFO for San Francisco, and LA for Los Angeles). The remaining columns indicate whether the SGO-reported crash belonged to each of the groups analyzed in this paper and described in the methodology section. As noted in the Methods section, two crashes occurred before the NHTSA SGO reporting requirements were enacted but were retroactively reviewed and deemed to meet the SGO reporting requirements (events #1 and #4 from Victor et al. (2023).
Police-reported status was determined using a combination of the NHTSA SGO field “Law Enforcement Investigating?” and internal Waymo records. Of the 10 SGO-reported crashes that indicated “Yes” for the “Law Enforcement Investigating?” field. The two SGO-reported crashes with an “Unknown” reported in the “Law Enforcement Investigating?” field also likely had a police report generated based on review of Waymo operations records. Additionally, in two events (30270-6548 and 30270-1583), the SGO field “Law Enforcement Investigating?” is “No”, but operational records indicate a police report may have been filed. These two events are also conservatively included in the SGO police-reported category. Event #1 from Victor et al. (2023) occurred before the NHTSA SGO-reporting period but was reported to police based on review of operational records.
@@ Appendix A Appendix
30270-6668
NA
PHX
True
True
True
False
False
30270-6667
NA
SFO
True
True
False
False
False
30270-$6$
NA
SFO
True
True
False
False
False
30270-6664
NA
PHX
True
True
False
False
False
30270-6663
NA
SFO
True
True
True
False
False
30270-6662
NA
SFO
True
True
False
False
False
30270-6661
NA
SFO
True
True
True
False
False
30270-6566
NA
PHX
True
True
False
True
True
30270-6561
NA
PHX
True
True
False
True
False
30270-6548
NA
PHX
True
True
False
True
False
30270-6542
NA
PHX
True
False
False
False
False
30270-6521
NA
PHX
True
True
False
False
False
30270-6520
NA
PHX
True
True
False
False
False
30270-6519
NA
PHX
True
True
True
True
False
30270-6518
NA
SFO
True
False
False
False
False
30270-6517
NA
SFO
True
True
False
False
False
30270-6516
NA
PHX
True
True
True
False
False
30270-6515
NA
SFO
True
True
True
False
False
30270-6514
NA
SFO
True
True
False
False
False
30270-6405
NA
SFO
True
True
True
True
False
30270-6385
NA
SFO
True
True
True
False
False
30270-6368
NA
PHX
True
True
True
True
False
30270-6352
NA
SFO
True
True
True
True
False
30270-6347
NA
LA
True
True
True
False
False
30270-6346
NA
SFO
True
True
False
False
False
30270-6342
NA
SFO
True
True
False
False
False
30270-6338
NA
PHX
True
True
True
False
False
30270-6337
NA
PHX
True
False
False
False
False
30270-6336
NA
SFO
True
True
True
True
True
30270-6335
NA
PHX
True
False
False
False
False
30270-6198
NA
PHX
True
True
True
True
False
30270-6150
NA
PHX
True
True
True
False
False
30270-6149
NA
PHX
True
True
True
True
False
30270-6133
NA
SFO
True
True
False
False
False
30270-6132
NA
SFO
True
True
False
False
False
30270-6131
NA
PHX
True
False
False
False
False
30270-5997
NA
PHX
True
True
True
True
True
30270-5961
NA
SFO
True
True
False
False
False
30270-5960
NA
SFO
True
False
False
False
False
30270-5959
NA
SFO
True
True
True
False
False
30270-5958
NA
SFO
True
False
False
False
False
30270-5957
NA
SFO
True
True
False
False
False
30270-5896
NA
PHX
True
True
True
True
False
30270-5760
NA
PHX
True
True
False
False
False
30270-5758
NA
SFO
True
True
True
False
False
30270-5756
NA
PHX
True
True
True
False
False
30270-5612
NA
PHX
True
True
False
False
False
30270-5611
NA
PHX
True
True
False
False
False
30270-5593
NA
SFO
True
False
False
False
False
30270-5592
NA
SFO
True
True
True
False
False
30270-5591
NA
PHX
True
True
True
False
False
30270-5588
NA
PHX
True
True
False
False
False
30270-5456
NA
SFO
True
True
True
False
False
30270-5319
NA
PHX
True
True
False
False
False
30270-5318
NA
SFO
True
True
False
False
False
30270-5208
NA
SFO
True
True
False
False
False
30270-5114
NA
SFO
True
True
False
False
False
30270-5085
NA
SFO
True
True
True
False
False
30270-5083
NA
SFO
True
True
True
False
False
30270-5081
NA
SFO
True
True
False
False
False
30270-4882
NA
PHX
True
True
True
False
False
30270-4880

### v2

@@ DISCUSSION
Although the benchmarks used in this study could adjust for some factors that are related to crashed vehicle rate (e.g., geographical area, types of roads, vehicle type), other pertinent factors were not controlled for. These include time of day adjustments and adjustments for the driving density within the geographic areas. As the Waymo ADS is being used as a ride-hailing service, the routes the ADS drives are likely differently dispersed than the general driving population and may be biased toward more densely populated areas. There are challenges when matching benchmark crash and VMT data sources that can be adjusted by these parameters, particularly in VMT data. An aggregate-level analysis comparing the ADS to the current crash status quo should include the ADS route choices, even if different from comparable human drivers, if the research question (like the one posed in this paper) is to determine the safety impact of replacing human driving with ADS vehicle driving.
An ADS vehicle operating in a ride hailing context is not always occupied by a human (e.g., before or in between picking up passengers). Therefore, when examining injury outcomes, if the ADS vehicle is unoccupied, there would be a reduced risk of injury for a similar human-only crash. As in the route choice discussion above, the objective of this study was to compare the ADS crash involvement to current human drivers, so it is appropriate to include miles where the ADS vehicle is not occupied. More broadly, risk of injury in a motor vehicle crash can be affected by many factors, such as seat belt use and vehicle model year (as a proxy for available safety systems). A risk-based approach that uses an injury risk function to compute probability of an injury outcome regardless of the whether an occupant is in the vehicle given some set of inputs can be advantageous in normalizing risk in both the benchmark and ADS crashes for comparison, as has been traditionally done in prospective safety benefit analyses (for example, Kusano and Gabler (2012); Scanlon et al. (2017)).
The Waymo RO service studied in this paper was operating over multiple years on different vehicle platforms and with different software versions. Just as the human driving population has changing characteristics over time, so do certain aspects of the Waymo RO ride-hailing service. The operating environment and territory have expanded during this time period. Starting in the area around Chandler, Arizona in 2020, today the Waymo RO ride-hailing service operates in an area that includes multiple cities in metropolitan Phoenix and over almost the entire area of San Francisco. A limitation of this study is that the entire history of the Waymo RO crash record was compared to human crash data over different time periods. As ADS fleets continue to increase the number of VMT, additional analyses that are restricted to certain periods of time or platforms could be possible. Many past retrospective studies of safety technologies, however, group different implementations of the same technology over long time periods to investigate the overall effect of a technology. For example, vehicles from different manufacturers were compared to investigate the effectiveness of first- and second-generation airbags Olson et al. (2006) and electronic stability control Riexinger et al. (2019). Thus, aggregate performance of RO ADS are still informative at this point in time.
CONCLUSIONS
This study compared the crashed vehicle rate of the Waymo RO service to human benchmarks for any property damage or injury, police-reported, and any-injury-reported outcomes. The results show that the Waymo RO service had a lower police-reported and any-injury-reported crashed vehicle rate compared to the benchmarks. When considering all locations together, the any-injury-reported crashed vehicle rate was 0.41 incidents per million miles (IPMM) for the ADS vs 2.80 IPMM for the human benchmark, an 85% reduction or a human crash rate that is 6.7 times higher than the ADS rate. Police-reported crashed vehicle rates for all locations together were 2.1 IPMM for the ADS vs. 4.68 IPMM for the human benchmark, a 55% reduction or a human crash rate that was 2.2 times higher than the ADS rate. Because the police-reported and any-injury-reported crashed ADS vehicle rates were much lower than the benchmark rates, the ADS any-injury-reported and police-reported crashed vehicle rates were statistically significant over a total of 7.14 million RO miles when compared individually in Phoenix and San Francisco and when compared overall. Although these results are promising, they are based on a small number of ADS events (15 police-reported and 3 any injury-reported). Thus, these comparisons may be subject to statistical noise as more miles and events are observed in the future. The comparison for Los Angeles was not statistically significant for zero observed ADS events but few miles. The results for the police-reported and any-injury-reported outcome groups, however, show a positive safety impact of the ADS in these types of crashes.
The results are more difficult to interpret for the any property damage or injury outcome comparison. The direction of the result and the statistical significance of the comparison depended on which benchmark was used (under-reporting adjusted police-report or NDS data) and whether low delta-V (< 1 mph) crashes were excluded from the ADS comparison. The results of the any property damage or injury comparison showed a statistically significant decrease in crashes in Phoenix when the ADS was compared to a police-reported benchmark adjusted for underreporting and in San Francisco and overall when compared to NDS benchmarks. There was an increase in the ADS crash rate when compared to an underreporting adjusted benchmark in San Francisco that was not statistically significant. Half of all ADS crashes had a delta-V less than 1 mph. When excluding ADS crashes with a delta-V less than 1 mph, the decrease in any property damage or injury crash rate for the ADS was statistically significant in all comparisons except in the comparison to the underreporting adjusted benchmark from San Francisco. These inconclusive results are largely due to an unclear lower reporting threshold in the human benchmarks. More research is needed to develop benchmarks that have an objective lower reporting threshold to draw conclusions about the any property damage or injury outcome group.
This is one of the first studies to compare overall crashed vehicle rates of ADS RO data only (as opposed to a mix of ADS testing with a human behind the wheel) to human benchmarks that also corrected for biases reported in the literature. The comparison of crash rates to benchmarks is part of the safety determination lifecycle, which itself is a part of a broader ADS safety case. The results show that the Waymo ADS operating in RO configuration has a lower crashed vehicle rate than human drivers when considering police-reported and any-injury-reported crashes. This result provides directional confirmation of the predictions and analyses used prior to deploying RO operations.
DISCLOSURE STATEMENT
All authors are employed by Waymo LLC.
Appendix A Appendix
A.1 Selection of SGO Events
The table below lists all the crashes included in this analysis. The SGO report ID can be used to look up further details of these events that are published by NHTSA. Detailed narratives or other data is not included in this paper, as this information can be obtained from the NHTSA SGO reporting website (https://www.nhtsa.gov/laws-regulations/standing-general-order-crash-reporting). The “Waymo 1M Mile RO Event #” column lists the integer event identifier that was reported in Victor et al. (2023) that described all contact events (regardless of whether they were reported as part of the NHTSA SGO). Of the 20 events reported in Victor et al. (2023), 11 were reported as part of the NHTSA SGO. Waymo started to operate in RO before the NHTSA SGO was enacted in July 2021 and as a result, 4 contact incidents that were previously reported in Victor et al. (2023) were not subject to the SGO reporting requirements. The authors retroactively reviewed these 4 contact incidents and determined that 2 of the contact incidents would have qualified for NHTSA SGO reporting. These two (2) incidents were included in the data analyzed for this paper. The location column refers to the market the crash occurred in (PHX for Phoenix, SFO for San Francisco, and LA for Los Angeles). The remaining columns indicate whether the SGO-reported crash belonged to each of the groups analyzed in this paper and described in the methodology section. As noted in the Methods section, two crashes occurred before the NHTSA SGO reporting requirements were enacted but were retroactively reviewed and deemed to meet the SGO reporting requirements (events #1 and #4 from Victor et al. (2023).
In the SGO data, crashes reported by Waymo were identified using the “Reporting Entity” field having the value “Waymo LLC”. The field “Driver / Operator Type” was set to “None” for RO operations, and set to “In-Vehicle (Commercial / Test)” when an autonomous specialist is monitoring the ADS behind the steering wheel. Note that the 7.1 million miles of ADS vehicle driving that were used to compute the rates presented in the paper only include RO miles. These 7.1 million RO miles should not be used to compute rates for SGO-reported crashes where an autonomous specialist was behind the wheel. The NHTSA SGO requires ADS fleet operators to report crashes where the ADS vehicle contributed or is alleged to have contributed (by steering, braking, acceleration, or other operational performance) to another vehicle’s physical impact with another road user or property involved in a reportable crash even if the ADS vehicle was not involved in the physical impact. Two Waymo RO collisions during the study period involved alleged contribution but no damage to the Waymo vehicle. These two cases are described in more detail in the appendix.
An administrative error in Waymo’s SGO submissions prior to February 2022 resulted in some duplicate crash entries in the published NHTSA SGO data. The NHTSA SGO has a facility for updating or making corrections to previously submitted crashes. An updated submission may be made if new information is available that was not available at the time the original report was submitted. Prior to February 2022, Waymo submitted a new SGO report for some crashes that were intended to be updates to previous crashes. This resulted in a new SGO report ID being issued instead of updating the information for the original crash. The event narratives in the published SGO data for these updated reports with new report IDs state they are duplicate submissions that update some information, and reference the original SGO report ID. The following NHTSA report IDs were excluded from this analysis because they are duplicates of updated reports: 30270-2248, 30270-2198, 30270-2168, 30270-2160, 30270-1949, 30270-1815, 30270-1160, 30270-1613, 30270-1748, 30270-1778, 30270-1535.
An “in-transport” vehicle is one that is traveling (moving or stopped) in the roadway. “Not-in-transport” vehicles are those that are parked in a designated parking spot out of the roadway. Therefore, including ADS vehicle crashes where the ADS vehicle is a not in-transport vehicle would inflate the ADS crash count relative to the benchmark. The ADS vehicle was considered not in-transport if the ADS vehicle was in the “park” gear and the ADS vehicle was in a designated parking area. This not in-transport status can happen when the ADS vehicle is parked awaiting its next trip instructions. Because the ADS software is running and in control of the vehicle, contacts while the ADS vehicle is not in-transport are reported as part of the NHTSA SGO reporting requirements. If the parking spot was on-street parking parallel to the traffic flow, the ADS vehicle must have been within 18 inches from the curb of the road as measured by the measured sensor data and curb location in the ADS map to be considered not-in-transport. Instances where the ADS vehicle was in “park” gear, but not near the curb were included as in-transport vehicles. The alternative, to include “not-in-transport” vehicle ADS crashes, would require also including “not-in-transport” vehicles in the benchmark, which are included in separate counts from in-transport vehicles in most police report databases, if at all. Additionally, it is unclear how a rate of crashed vehicles per VMT should be interpreted for not-in-transport vehicles, as parked vehicle crash risk is likely more related to amount of time parked rather than VMT. Underreporting for not-in-transport vehicles may also differ from in-transport vehicles, which was not accounted for the underreporting adjustments used in the benchmarks (Scanlon et al., 2023; Blincoe et al., 2023).
The “in-transport” determination was made with ADS sensor and map data that are not part of the crash data published as part of the SGO. The event narratives in the NHTSA SGO data for Waymo can also be used to determine whether the vehicle was parked in a valid parking spot. There were 12 SGO crashes where the Waymo vehicle was in “park” gear at the time of the crash. Of those 12, in 7 crashes the Waymo vehile was parked near the curb and in 5 the Waymo vehicle was not near the curb. In 6 out 7 crashes where the Waymo vehicle was near the curb, the event narratives state the vehicle was parked “near” or “adjacent” to the curb. Of the 8 crashes where the Waymo vehicle was in park but not in a parking spot or near the curb, 2 of the narratives mentioned the vehicle had pulled “near” or “pulled to” the curb. Therefore, using the event narratives alone would yield similar results than was found using the additional sensor data (8 in park and near the curb and 7 in park but not near a curb using narratives compared to 7 in park and near the curb and 8 in park but not near a curb when using sensor and map data). There is no reporting requirement in the SGO for parking status nor is there a definition of a “valid” parking space. Having agreed upon definitions for “in-transport” status would improve comparability between ADS crash data and human crash data.
Two NHTSA SGO-reported crashes did not include damage to the ADS vehicle. First, case 30270-6542 involved two human-driven vehicles that made contact while following behind a Waymo ADS vehicle. The Waymo vehicle was not contacted or damaged. Second, case 30270-6341 involved a Waymo ADS vehicle stopped near the entrance of an intersection yielding to a fire engine. While the fire engine was navigating around the stopped Waymo vehicle, the fire engine allegedly contacted another vehicle also stopped in a through lane. Because the purpose of this paper is to compare crashed vehicle rates between ADS and human benchmarks, we excluded these crashes from the NHTSA SGO in-transport vehicle category and subsequent categories. These two events were included in the SGO-reported count.
Police-reported status was determined using a combination of the NHTSA SGO field “Law Enforcement Investigating?” and internal Waymo records. There were 10 SGO-reported crashes that indicated “Yes” for the “Law Enforcement Investigating?” field. The two (2) SGO-reported crashes with an “Unknown” reported in the “Law Enforcement Investigating?” field also likely had a police report generated based on review of Waymo operations records. Additionally, in two events (30270-6548 and 30270-1583), the SGO field “Law Enforcement Investigating?” is “No”, but operational records indicate a police report may have been filed. These two events were also conservatively included in the SGO police-reported category. Event #1 from Victor et al. (2023) occurred before the NHTSA SGO-reporting period but was reported to police based on review of operational records. To determine any-injury-reported crashes, the “Highest Injury Severity Alleged” SGO field was used. Any crash that did not have “No Injuries Reported” was considered an any-injury-reported crash. In this dataset, there were 3 crashes with “Minor” injuries reported. Crashes where the ADS was not-in-transport were also excluded from police-reported and alleged any-injury-reported crashes. All ADS not-in-transport crashes had no reported law enforcement investigation and no injuries reported as part of the SGO data.
@@ A.1 Selection of SGO Events
30270-6668 NA PHX True True True False False
30270-6667 NA SFO True True False False False
30270-$6$ NA SFO True True False False False
30270-6664 NA PHX True True False False False
30270-6663 NA SFO True True True False False
30270-6662 NA SFO True True False False False
30270-6661 NA SFO True True True False False
30270-6566 NA PHX True True False True True
30270-6561 NA PHX True True False True False
30270-6548 NA PHX True True False True False
30270-6542 NA PHX True False False False False
30270-6521 NA PHX True True False False False
30270-6520 NA PHX True True False False False
30270-6519 NA PHX True True True True False
30270-6518 NA SFO True False False False False
30270-6517 NA SFO True True False False False
30270-6516 NA PHX True True True False False
30270-6515 NA SFO True True True False False
30270-6514 NA SFO True True False False False
30270-6405 NA SFO True True True True False
30270-6385 NA SFO True True True False False
30270-6368 NA PHX True True True True False
30270-6352 NA SFO True True True True False
30270-6347 NA LA True True True False False
30270-6346 NA SFO True True False False False
30270-6342 NA SFO True True False False False
30270-6341 SFO True False False False False
30270-6338 NA PHX True True True False False
30270-6337 NA PHX True False False False False
30270-6336 NA SFO True True True True True
30270-6335 NA PHX True False False False False
30270-6198 NA PHX True True True True False
30270-6150 NA PHX True True True False False
30270-6149 NA PHX True True True True False
30270-6133 NA SFO True True False False False
30270-6132 NA SFO True True False False False
30270-6131 NA PHX True False False False False
30270-5997 NA PHX True True True True True
30270-5961 NA SFO True True False False False
30270-5960 NA SFO True False False False False
30270-5959 NA SFO True True True False False
30270-5958 NA SFO True False False False False
30270-5957 NA SFO True True False False False
30270-5896 NA PHX True True True True False
30270-5760 NA PHX True True False False False
30270-5758 NA SFO True True True False False
30270-5756 NA PHX True True False False False
30270-5612 NA PHX True True False False False
30270-5611 NA PHX True True False False False
30270-5593 NA SFO True False False False False
30270-5592 NA SFO True True True False False
30270-5591 NA PHX True True True False False
30270-5588 NA PHX True True False False False
30270-5456 NA SFO True True True False False
30270-5319 NA PHX True True False False False
30270-5318 NA SFO True True False False False
30270-5208 NA SFO True True False False False
30270-5114 NA SFO True True False False False
30270-5085 NA SFO True True True False False
30270-5083 NA SFO True True True False False
30270-5081 NA SFO True True False False False
30270-4882 NA PHX True True True False False
30270-4880 19 PHX True True False False False
30270-4768 20 PHX True True True False False
30270-4484 16 PHX True True False True False
30270-4363 17 PHX True True False False False
30270-3842 14 SFO True True True False False
30270-3838 13 PHX True True False False False
30270-1730 6 PHX True False False False False
30270-1583 7 PHX True True True True False
30270-1501 5 PHX True True True False False
NA 4 PHX True True True False False
NA 1 PHX True True True True False
As stated in the Methods section (section METHODS), the “NHTSA SGO In-Transport Exclude Low delta-V (DV)” category was derived by applying a impulse-momentum collision model for vehicle-to-vehicle crashes. The collision model used was an impulse-momentum model using the inertial properties of the ADS equipped vehicle and estimated inertial properties of the crash partner, as described in Scanlon et al. (2021). Each of the 7 crashes with fixed or non-fixed objects was examined individually to estimate a delta-V. Of the 7 crashes with fixed or non-fixed objects, 5 were excluded for having a low delta-V. Crashes where the ADS was not-in-transport were also excluded from this group. Three (3) fixed object crashes resulted in underbody damage when the Waymo vehicle “bottomed out” entering a parking lot driveway (30270-5612, 30270-5611) or tire damage after driving through a pothole (30270-5114) which resulted in less than the 1.0 mph delta-V threshold. Two (2) crashes involved non-fixed objects with low mass (<100 kg): a cardboard box (30270-5081) and a swinging parking lot gate that was not fixed (30270-4363). The single crash involving a cyclist (30270-5456) was included in this group.
Of the two vehicle-to-object crashes included, SGO report 30270-6548 involved a Waymo ADS vehicle that was driving in a construction zone and “entered a lane undergoing construction …, encountered a section of roadway that had been removed, and the front driver’s side wheel dropped off the paved roadway.” After reviewing the on-board data from this crash event, the front wheel dropping likely caused a delta-V of 1 to 3 mph. SGO report 30270-6561 involved a Waymo ADS vehicle “exited a parking lot…, when the passenger’s side sensors and side mirror made contact with an automatic gate that was in the process of closing.” Based on review of the sensor data, the contact with the closing gate caused damage to a sensor housing and side view mirror of the vehicle, which are unlikely to result in a delta-V above 1 mph. Because the event in report 30270-6561 was police-reported, we decided to include this event from the low delta-V category.
A.2 Results for Los Angeles
In Los Angeles, there was a single SGO-Reported crash during the study period. This single event also met the in-transport and exclude low delta-V inclusion criteria resulting in an IPMM of 21.4 with 95% confidence intervals of (0.5, 119) IPMM. There were no (zero) police-reported and any-injury-reported crashes in Los Angeles, resulting in a 0 IPMM with 95% confidence intervals of (0, 64.1) IPMM.
The 1 crash in Los Angeles met both the SGO and low delta-V thresholds in the any property damage or injury comparison. Compared to the Blincoe-adjusted benchmark, the rate ratio was 2.84 with confidence intervals of (0.04, 18.1). Compared to the Blanco-adjusted benchmark, the rate ratio was 1.40 with confidence intervals of (0.02, 8.92). Due to the low ADS miles driven, neither comparison was statistically significant.
There were no observed police-reported or any-injury-reported crashes in Los Angeles resulting in an rate ratio of 0. The result in Los Angeles was not statistically significant due to low mileage (confidence intervals of [0, 21.3] for police-reported and [0, 32.8] for any-injury-reported).
A.3 Example of Nelson (1970) Rate Ratio Confidence Intervals
An example of the Nelson (1970) confidence interval for the crash rate ratio is shown in Listing 1 in the Python language.
Listing 1: Python Function to Compute Nelson (1970) Rate Ratio Confidence Intervals
1from scipy.stats import betaprime
2
3def nelson_ci(X, s, Y, t, alpha=0.05):
4 """Clopper-Pearson limits for a rate ratio.
5
6 Args:
7 X: Benchmark event count
8 s: Benchmark exposure (e.g., miles)
9 Y: ADS Event Count
10 t: ADS exposure (e.g., miles, matching benchmark exposure)
11 alpha: alpha level
12
13 Returns:
14 Rate Ratio and (lower, upper) confidence intervals
15 """
16 if Y > 1e-3:
17 # 2-sided test, except for zero events (then it’s one sided)
18 alpha = alpha / 2